OpenStudy (anonymous):

A quantity that depends on force, distance, and time is called ____. A.momentum B.work C.displacement D.power

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

momentum \(p=mv\) work \(W=Fd\) displacement \(d_2-d_1\) Power \(p=\frac{W}{t}=\frac{Fd}{t}\) so which one do you think?

OpenStudy (anonymous):

B. work?

OpenStudy (anonymous):

nope, work only deals with force and displacement

OpenStudy (anonymous):

no wait D. power

OpenStudy (anonymous):

A quantity that depends on force, distance, and time is called ____. d.power

OpenStudy (anonymous):

yes @Ballin_24

OpenStudy (abhisar):

@lupita1995 Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem. Seeking/Giving Just/Direct Answers is against the \(\large\sf\color{red}{\underline{\href{ /code-of-conduct}{CODE~OF~CONDUCT}}}\)

OpenStudy (anonymous):

sorry for helping

OpenStudy (anonymous):

no its ok thanks for helping me (: @lupita1995

OpenStudy (anonymous):

np